OpenStudy (anonymous):

z=x+(x+1)i the mod of z is 5. find x

OpenStudy (anonymous):

mod of z=a+bi=sqrt(a^2+b^2)

OpenStudy (anonymous):

is it 3 and -4

OpenStudy (anonymous):

nvm it is

OpenStudy (anonymous):

\[x^2+(x+1)^2=25\]\[2x^2+2x-24=0\]\[x^2+x-12=0\]\[(x+4)(x-3)=0\]u r right

OpenStudy (anonymous):

hm can you find the argument of Z from this ?

OpenStudy (anonymous):

arg1=atan(4/5) and arg2=-pi+atan(3/4)

OpenStudy (anonymous):

is that right??

OpenStudy (anonymous):

almost i got different values for arg1

OpenStudy (anonymous):

z=3+4i theta=Arctan(4/3)

OpenStudy (anonymous):

yeah oops thought 5 was 3.. wb the second arg

OpenStudy (anonymous):

i got the same for the 2nd